A lateral edge of the hexagonal pyramid is 2.0m long and its base perimeter is 6m determine the volume of the pyramid and its lateral area
1
Expert's answer
2020-03-10T11:16:05-0400

Volume of Hexagonal pyramid is given by "\dfrac{\sqrt3}{2}a^2h" ; where a is the base edge and h is the height

base perimeter = 6*a

6=6*a

a=1 m

lateral edge ="\sqrt{a^2+h^2}" "=\sqrt{1^2+h^2}=2,"

"h=\sqrt{3}"

Volume = "\dfrac{\sqrt3}{2}a^2h" ="\dfrac{\sqrt3}{2}1^2\sqrt3=3/2"


Lateral surface area = "3a\sqrt{h^2+\dfrac{3a^2}{4}}=3*1*\sqrt{3+3/4}=3\sqrt{15}/2"